Project

General

Profile

Milestone #25216

Testing and Installing 2 RCC at CMTF

Added by Philip Varghese 3 months ago. Updated about 1 month ago.

Status:
Resolved
Priority:
Normal
Category:
SSR1 Resonance Control
Start date:
11/16/2020
Due date:
% Done:

100%

Estimated time:
4.00 h
Spent time:
Duration:

Description

2 RCC chassis are to be configured , tested and installed at CMTF for
SSR1.
Network configuration, fiber link to LLRF controller testing are being done on one
RCC and the other is installed at CMTF. These chassis will be shared with the LCLS-II based
RFS controller to be tested with SSR1 and have entries in that project also. The information is also relevant
for the Meson 650STC project.

SSR1_RCC_Config_v2_Oct29_2020.docx (15.4 KB) SSR1_RCC_Config_v2_Oct29_2020.docx Philip Varghese, 11/16/2020 03:52 PM
SSR1_RCC_Config_Oct29_2020.docx (12.8 KB) SSR1_RCC_Config_Oct29_2020.docx Philip Varghese, 11/16/2020 03:52 PM
SSR1_Dev_I2C_Debug_Nov3_2020.docx (60.9 KB) SSR1_Dev_I2C_Debug_Nov3_2020.docx Philip Varghese, 11/16/2020 03:52 PM
SOCMFC2RCC_CommTest100HzWvfrm.PNG (147 KB) SOCMFC2RCC_CommTest100HzWvfrm.PNG Philip Varghese, 11/16/2020 03:56 PM
RCC_PIP_II_Test_SSR1.docx (180 KB) RCC_PIP_II_Test_SSR1.docx Test_listing Philip Varghese, 11/16/2020 03:58 PM
SSR1_RCC_CopyBitFile_Nov19_2020.docx (15.9 KB) SSR1_RCC_CopyBitFile_Nov19_2020.docx Philip Varghese, 11/19/2020 12:25 PM
SSR1_Controller1_To_RCC_CommTest100HzWvfrm.PNG (78 KB) SSR1_Controller1_To_RCC_CommTest100HzWvfrm.PNG SSR1 SocMFC1 to RCC comm link test Philip Varghese, 12/07/2020 04:22 PM
SSR1_Controller2_To_RCC_CommTest100HzWvfrm.PNG (145 KB) SSR1_Controller2_To_RCC_CommTest100HzWvfrm.PNG Philip Varghese, 12/09/2020 10:54 AM
SSr1_QSFP_Config2.png (113 KB) SSr1_QSFP_Config2.png Philip Varghese, 12/09/2020 10:57 AM
SSR1_Cavity1_Piezo_Disabled.png (159 KB) SSR1_Cavity1_Piezo_Disabled.png Philip Varghese, 12/15/2020 04:58 PM
SSR1_Cavity1_Piezo_Enabled.png (160 KB) SSR1_Cavity1_Piezo_Enabled.png Philip Varghese, 12/15/2020 04:58 PM
SSR1_Cavity1_Piezo_Control_Waveforms.PNG (324 KB) SSR1_Cavity1_Piezo_Control_Waveforms.PNG Philip Varghese, 12/15/2020 04:58 PM
SSR1_Cavity1_Piezo_Control_Settings.PNG (279 KB) SSR1_Cavity1_Piezo_Control_Settings.PNG Philip Varghese, 12/15/2020 04:58 PM
SSR1_Cavity1_Piezo_Control_LinkStatus.PNG (256 KB) SSR1_Cavity1_Piezo_Control_LinkStatus.PNG Philip Varghese, 12/15/2020 04:58 PM

History

#1 Updated by Philip Varghese 3 months ago

The test procedure outlined in the document RCC@PIP-II was exercised on the SOCMFC - RCC setup in the lab.
A test waveform was successfully received and plotted as shown in the attached transcipt.

#3 Updated by Philip Varghese 3 months ago

  • Estimated time deleted (4.00 h)

#4 Updated by Philip Varghese 3 months ago

The FPGA configuration bit file update procedure was exercised successfully.
The listing of commands and messages is attached.

#5 Updated by Philip Varghese 3 months ago

Controller 1 to the RCC fiber link has been successfully tested with a test waveform.

#6 Updated by Philip Varghese 3 months ago

The second SSR1 controller for cavity 3,4 was connected to the RCC and the link was tested
with the test waveform from the socmfc. The data was received successfully at the RCC.
Although there are three QSFP connectors on each RCC, Sergio requested that both connections
from the two socmfc controllers be made to a single QSFP connector on the RCC. This configuration
is shown in the attached diagram. Some patch cables and connectors are needed for this purpose- they
have been ordered and should be here by tomorrow.

#7 Updated by Philip Varghese 2 months ago

After a lot of work put in by the team, the first SSR1 cavity piezo tuner was successfully tested in closed loop.
It was nice to see the whole mix of hardware and software from the fermilab LLRF controller to the LCLS-II RCC
work with ACNET/LABVIEW/EPICS to run the piezo tuner system.
The results are attached here for those who cannot access the e-log (182817) entry.

#8 Updated by Shrividhyaa Sankar Raman about 1 month ago

  • % Done changed from 50 to 100
  • Status changed from New to Resolved

Also available in: Atom PDF